A 105-inch board is cut into two pieces. One piece is four times the length of the other. Find the lengths of the two pieces.
Answer by ikleyn(52788)

.
x + 4x = 105


5x = 105


x = 105/5 = 21.


The pieces are 21 and 4*21 = 84 inches.    ANSWER